Question 3

Essendon Pty Limited manufactures guidance systems for rockets that are used to launch

commercial satellites. The company’s Software Business Division reported the following data

for the last year:

Sales revenue $2,000,000

Sales returns $110,000

Cost of goods sold $1,200,000

Operating expenses $550,000

Total assets at year-beginning $1,900,000

Total assets at year-end $1,990,000

Total current liabilities at year-beginning $55,000

Total current liabilities at year-end $61,000

The company’s required rate of return is 8.75 percent. Additional information:

• The total assets at year-end include a piece of vacant land valued at $320,000. It is

identified as a non-productive asset by the corporate management.

• The divisional manager manages all current liabilities.

• The use of average balances is recommended.

In an attempt to improve its return on investment (ROI), Essendon company is planning to:

• speed-up the collection of all account receivables by $70,000.

• write-off and discard $57,000 of obsolete inventory.

• sponsor a competition for teams of local universities’ students to design and build

small satellite with budget up to $120,000

Required:

a. Calculate the Software Business Division’s return on investment (ROI) and residual

income (RI). Discuss the results from your calculations.

Note: Textbook Chapter 13 “Learning objective 13.5-Measuring Profit and Invested

capital” will be a good guide to define profit and invested capital in the ROI and RI

calculations.

b. Explain should the above plans be adopted in order to improve its ROI for Software

Business Division?

c. If profit and sales remain the same in the coming year, but the investment turnover

increases to 0.92, calculate the new ROI?
Answer #1

A. CAlculation of ROI

Particulars Amount
Sales 2000000
less:
Sales return 110000
Cost of goods sold 1200000
operating expenses 550000
Residual Income 140000
Average assets (1900000+1990000)/2 1945000
Average current liabilities (55000+61000)/2 58000
Net investments (1945000-58000) 1887000
Return on investment (140000/1887000) 7.42%

The performance of the company is not upto mark. The expected return is 8.75% but actually it is nearby 7.5%. The compnay needs to improve in terms of its sales. Also its high time to refine its assets and scap the non using assets.

2. The adoption of above plans will not make any material dofference. Because the collection of receivanbles will not make any change in total assets. Similarly development costs are capitalised and created as new asset. So it will also not make any change. If land is scrapped then it will obviously make a huge change but nothing is mentioned in above sum about this.

Calculation of revised net assets

Particulars Amount
Average net assets 1887000
less: obsolete inventory 57000
New net assets 1830000
Return on investments (140000/1830000) 7.65%

3. Investment turnover ratio =net sales/ net investments

                         0.92             =(2000000-110000)/ net investments

Net investments = 1890000/0.92 =2054347

New ROI = 140000/2054347 =6.81%